# Scalable Decentralized Algorithms for Online Personalized Mean Estimation

### v3

@@ Appendix B Appendix - Schematic Representation for B-ColME
Appendix C Appendix - Proof for B-ColME - Theorem 4
@@
Here we provide a convergence result in the sense of the above definition for B-ColME, we will derive a similar result for C-ColME in Appendix D.
@@
$\mathbb{P}\left(\forall t>\tau_{a}^{\prime},|\hat{\mu}_{a}^{t}-\mu_{a}|<\varepsilon\right)\geq 1-\delta$
$\qquad\text{with}\quad\tau_{a}^{\prime}=\max\left[\zeta_{D}+d,\frac{\widetilde{n}_{\frac{\delta}{2}}}{|\mathcal{CC}^{d}_{a}|}+d\right]$
@@
indeed observe that $\bar{x}(n)$, as immediate consequence of the elementary properties of normal random variables, is distributed as a Gaussian, with zero mean and standard deviation equal to $\frac{\sigma}{\sqrt{n}}$.
@@
Now, to conclude the proof, observe that conditionally over the fact that $\mathcal{C}_{a}^{t}=\mathcal{C}_{a}\cap\mathcal{N}_{a}$ for every time $t\geq t_{0}-d$ , agent $a$ computes its average $\hat{\mu}_{a}$ as an average of samples collected by all agents in $\mathcal{CC}_{a}^{d}$. Such samples are i.i.d. and follows $D_{a}$; moreover its number easily lower bounded by $|\mathcal{CC}_{a}^{d}|(t_{0}-d)$. Therefore whenever $|\mathcal{CC}_{a}^{d}|(t_{0}-d)\geq\widetilde{n}_{\frac{\delta}{2}}$ by the definition of $\widetilde{n}_{\frac{\delta}{2}}$ we have
@@
Now previous approach rather immediately extends to the case in which $D_{a}$ are sub-Gaussian with parameter $\sigma$, since in this case $\bar{x}(n)=\frac{1}{n}\sum_{1}^{n}x_{t}$ of independent samples $x_{t}$ extracted from sub-Gaussian distributions with parameter $\sigma$ having the same average $\mu_{a}$ is sub-Gaussian with parameter $\sigma/\sqrt{n}$ and zero mean, and therefore by definition of sub-Gaussian (see Vershynin [2018] for a discussion about sub-Gaussian distributions and their properties):
@@
At last consider the case in which $\mu_{a}$ have a fourth central moment uniformly bounded bu $\mu_{4}$, and a variance uniformly bounded by $\sigma^{2}$. In this case considering the empirical $\bar{x}(n)=\frac{1}{n}\sum_{1}^{n}x_{t}$ of independent samples $x_{t}$ extracted from arbitrary distributions with the same average $\mu_{a}$, following the same approach as in the proof of Proposition 11 we can bound its fourth moment as:
@@
As last step, now we derive easy upper bounds for $\widetilde{n}_{\frac{\delta}{2}}$. We start from the last case. We have
@@
Now considering the Gaussian case, in this case we can exploit the following well known bounds for $Q(x)$:
@@
At last consider the sub-Gaussian case:
@@
Appendix D Appendix - Proofs for C-ColME - Theorem 6 and Theorem 7
It is convenient to consider an auxiliary system for which the consensus matrix can be arbitrary until time $\zeta_{d}$ and then switches to a situation where agents only communicate with their neighbours belonging to the same class, i.e., $W_{t}=W$ for any $t\geqslant\zeta_{d}$ and $W_{a,a^{\prime}}>0$ if and only if $a^{\prime}\in\mathcal{C}_{a}\cap\mathcal{N}_{a}$.
We will derive some bounds for the auxiliary system for any choice of the matrices $W_{1},W_{2},\dots W_{\gamma_{D}}$. With probability $1-\frac{\delta}{2}$ these bounds apply also to the original system under study, because with such probability each agent correctly detects the neighbours in the same class for any $t>\zeta_{D}$. From now on, we refer then to the auxiliary system.
D.1 Preliminaries
After time $\zeta_{D}$ the original graph has been then split in $C$ connected components, where component $c$ includes $n_{c}$ agents. By an opportune permutation of the agents, we can write the matrix $W$ as follows
@@ D.1 Preliminaries
For $t>\zeta_{D}$, the estimates in component $c$ evolves independently from the other components and we can write:
@@ D.2 Technical Results
where $e_{n}(t)\triangleq(n+1)!te^{-t}\int_{-\infty}^{t}\frac{e^{x}}{x^{n+2}}\mathop{}\!\mathrm{d}x$ for $n=0$. The remainder $e_{n}(t)$ can be bounded by the $n+1$-th term times the factor $1+\sqrt{\pi}\frac{\Gamma(n/2+3/2)}{\Gamma(n/2+1)}$ Digital Library of Mathematical Functions [2023]. This factor is smaller than $3$ for $n=0$.
@@
We will omit the indices when they run from $1$ to $t$ and denote by $\mathbf{z}$ and $\mathbf{z}^{\prime}$ two generic indipendent random variables distributed as $\mathbf{z}_{\tau}$.
@@
$=\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2},t_{3},t_{4}}\mathbbm{1}_{t_{1}\leqslant\tau_{1}}\mathbbm{1}_{t_{2}\leqslant\tau_{2}}\mathbbm{1}_{t_{3}\leqslant\tau_{3}}\mathbbm{1}_{t_{4}\leqslant\tau_{4}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]$ (43)
@@
$=\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2},t_{3},t_{4}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{3}\leqslant\tau_{3},t_{4}\leqslant\tau_{4}}\times\Big{(}\mathbbm{1}_{t_{1}=t_{2}=t_{3}=t_{4}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]$
$\phantom{=\sum}+\mathbbm{1}_{t_{1}=t_{2},t_{3}=t_{4},t_{1}\neq t_{3}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]$
$\phantom{=\sum}+\mathbbm{1}_{t_{1}=t_{3},t_{2}=t_{4},t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]$
${\phantom{=\sum}+\mathbbm{1}_{t_{1}=t_{4},t_{2}=t_{3},t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]\Big{)}}$ (44)
$=\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{1}\leqslant\tau_{2},t_{1}\leqslant\tau_{3},t_{1}\leqslant\tau_{4}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{1}}\mathbf{z}_{t_{1}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{1}}\right]$
@@
$\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{3}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{1}\leqslant\tau_{2},t_{3}\leqslant\tau_{3},t_{3}\leqslant\tau_{4}}\mathbbm{1}_{t_{1}\neq t_{3}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{1}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{3}}\right]$
@@
$\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{1}\leqslant\tau_{3},t_{2}\leqslant\tau_{4}}\mathbbm{1}_{t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{1}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{2}}\right]$
@@
${\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{2}\leqslant\tau_{3},t_{1}\leqslant\tau_{4}}\mathbbm{1}_{t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{2}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{1}}\right]}$ (45)
$=\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{1}\leqslant\tau_{2},t_{1}\leqslant\tau_{3},t_{1}\leqslant\tau_{4}}\operatorname*{\mathbb{E}}\left[\mathbf{z}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}\mathbf{z}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}\right]$
@@
$\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{3}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{1}\leqslant\tau_{2},t_{3}\leqslant\tau_{3},t_{3}\leqslant\tau_{4}}\mathbbm{1}_{t_{1}\neq t_{3}}\operatorname*{\mathbb{E}}\left[\mathbf{z}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}\right]\operatorname*{\mathbb{E}}\left[{\mathbf{z}^{\prime}}^{\top}A(\tau_{3},\tau_{4}){\mathbf{z}^{\prime}}\right]$
@@
$\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{1}\leqslant\tau_{3},t_{2}\leqslant\tau_{4}}\mathbbm{1}_{t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}^{\prime}\mathbf{z}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}^{\prime}\right]$
@@
${\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2}}\mathbbm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{2}\leqslant\tau_{3},t_{1}\leqslant\tau_{4}}\mathbbm{1}_{t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}^{\prime}\mathbf{z}^{\prime\top}A(\tau_{3},\tau_{4})\mathbf{z}\right]}$ (46)
@@
As $W$ is symmetric and stochastic, then the module of its largest eigenvalue is equal to $1$. The vector $\mathbf{1}_{n}$ is both a left and right eigenvector of $W$ relative to the simple eigenvalue $1$. Then, $P$ is the projector onto the null space of $W-I$ along the range of $W-I$ Meyer [2001][p.518]. The spectral theorem leads us to conclude that the eigenvalues of $W-P$ (counted with their multiplicity) are then all eigenvalues of $W$ except $1$ and with the addition of 00. We can then conclude that $\lVert W-P\rVert=\rho(W-P)=\lambda_{2}(W)$. $W$ is irreducible with non-negative elements on the diagonal, then it is primitive Meyer [2001][Example 8.3.3], i.e., $1$ is the only eigenvalue on the unit circle.
@@
where (57) follows from Lemma A.1 in Shalev-Shwartz and Ben-David [2014]. For $a<1$, it is easy to check that for $x\geqslant 1$, $x\geqslant\frac{a}{4}\ln^{2}x$ holds unconditionally. ∎
D.3 Bounding the $4$-th Moment of the Estimation Error.
@@
D.4 $(\varepsilon,\delta)$-Bounds: Proof of Theorem 7

### v4

@@ Appendix D Appendix D - Schematic Representation for B-ColME
Appendix E Appendix E - Proof for B-ColME - Theorem 4
@@
Here we provide a convergence result in the sense of the above definition for B-ColME, we will derive a similar result for C-ColME in Appendix F.
@@
$\mathbb{P}\left(\forall t>\tau^{B}_{a},|\hat{\mu}_{a}^{t}-\mu_{a}|<\varepsilon\right)\geq 1-\delta$
$\qquad\text{with}\quad\tau^{B}_{a}=\max\left[\zeta_{D}+d,\frac{\widetilde{n}_{\frac{\delta}{2}}}{|\mathcal{CC}^{d}_{a}|}+d\right]$
@@
indeed observe that $\bar{x}(n)$, as an immediate consequence of the elementary properties of normal random variables, is distributed as a Gaussian, with zero mean and standard deviation equal to $\frac{\sigma}{\sqrt{n}}$.
@@
Now, to conclude the proof, observe that conditionally over the fact that $\mathcal{C}_{a}^{t}=\mathcal{C}_{a}\cap\mathcal{N}_{a}$ for every time $t\geq t_{0}-d$, agent $a$ computes its average $\hat{\mu}_{a}$ as an average of samples collected by all agents in $\mathcal{CC}_{a}^{d}$. Such samples are i.i.d. and follow $D_{a}$; moreover, their number is easily lower bounded by $|\mathcal{CC}_{a}^{d}|(t_{0}-d)$. Therefore whenever $|\mathcal{CC}_{a}^{d}|(t_{0}-d)\geq\widetilde{n}_{\frac{\delta}{2}}$ by the definition of $\widetilde{n}_{\frac{\delta}{2}}$ we have
@@
Now, the previous approach rather immediately extends to the case in which $D_{a}$ are sub-Gaussian with parameter $\sigma$, since in this case $\bar{x}(n)=\frac{1}{n}\sum_{1}^{n}x_{t}$ of independent samples $x_{t}$ extracted from sub-Gaussian distributions with parameter $\sigma$ having the same average $\mu_{a}$ is sub-Gaussian with parameter $\sigma/\sqrt{n}$ and zero mean, and therefore by definition of sub-Gaussian (see (Vershynin 2018) for a discussion about sub-Gaussian distributions and their properties):
@@
At last, consider the case in which $\mu_{a}$ has a fourth central moment uniformly bounded by $\mu_{4}$, and a variance uniformly bounded by $\sigma^{2}$. In this case considering the empirical $\bar{x}(n)=\frac{1}{n}\sum_{1}^{n}x_{t}$ of independent samples $x_{t}$ extracted from arbitrary distributions with the same average $\mu_{a}$, following the same approach as in the proof of Proposition 11 we can bound its fourth moment as:
@@
As the last step, now we derive easy upper bounds for $\widetilde{n}_{\frac{\delta}{2}}$. We start from the last case. We have
@@
Now considering the Gaussian case, in this case, we can exploit the following well-known bounds for $Q(x)$:
@@
At last, consider the sub-Gaussian case:
@@
We observe that $\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)$ is smaller than the corresponding term $\beta_{\frac{\delta}{2}}^{-1}(\varepsilon)$ appearing in Theorem 8 for ColME. Our proofs can readily be adapted to ColME, enabling to substitute $n_{\frac{\delta}{2}}^{\star}(\varepsilon)$ in (Asadi et al. 2022) [Theorem 2] with the smaller term $\widetilde{n}_{\frac{\delta}{2}}(\varepsilon)$.
Appendix F Appendix F - Proofs for C-ColME - Theorem 2 and Theorem 3
It is convenient to consider an auxiliary system for which the consensus matrix can be arbitrary until time $\zeta_{d}$ and then switches to a situation where agents only communicate with their neighbors belonging to the same class, i.e., $W_{t}=W$ for any $t\geqslant\zeta_{d}$ and $W_{a,a^{\prime}}>0$ if and only if $a^{\prime}\in\mathcal{C}_{a}\cap\mathcal{N}_{a}$.
We will derive some bounds for the auxiliary system for any choice of the matrices $W_{1},W_{2},\dots W_{\gamma_{D}}$. With probability $1-\frac{\delta}{2}$ these bounds apply also to the original system under study because with such probability each agent correctly detects the neighbors in the same class for any $t>\zeta_{D}$. From now on, we refer then to the auxiliary system.
F.1 Preliminaries
After time $\zeta_{D}$ the original graph has been then split into $C$ connected components, where component $c$ includes $n_{c}$ agents. By an opportune permutation of the agents, we can write the matrix $W$ as follows
@@ F.1 Preliminaries
For $t>\zeta_{D}$, the estimates in component $c$ evolve independently from the other components and we can write:
@@ F.2 Technical Results
where $e_{n}(t)\triangleq(n+1)!te^{-t}\int_{-\infty}^{t}\frac{e^{x}}{x^{n+2}}\mathop{}\!\mathrm{d}x$ for $n=0$. The remainder $e_{n}(t)$ can be bounded by the $n+1$-th term times the factor $1+\sqrt{\pi}\frac{\Gamma(n/2+3/2)}{\Gamma(n/2+1)}$ (of Mathematical Functions 2023). This factor is smaller than $3$ for $n=0$.
@@
We will omit the indices when they run from $1$ to $t$ and denote by $\mathbf{z}$ and $\mathbf{z}^{\prime}$ two generic independent random variables distributed as $\mathbf{z}_{\tau}$.
@@
$=\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2},t_{3},t_{4}}\bm{1}_{t_{1}\leqslant\tau_{1}}\bm{1}_{t_{2}\leqslant\tau_{2}}\bm{1}_{t_{3}\leqslant\tau_{3}}\bm{1}_{t_{4}\leqslant\tau_{4}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]$ (43)
@@
$=\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2},t_{3},t_{4}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{3}\leqslant\tau_{3},t_{4}\leqslant\tau_{4}}\times\Big{(}\bm{1}_{t_{1}=t_{2}=t_{3}=t_{4}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]$
$\phantom{=\sum}+\bm{1}_{t_{1}=t_{2},t_{3}=t_{4},t_{1}\neq t_{3}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]$
$\phantom{=\sum}+\bm{1}_{t_{1}=t_{3},t_{2}=t_{4},t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]$
${\phantom{=\sum}+\bm{1}_{t_{1}=t_{4},t_{2}=t_{3},t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{4}}\right]\Big{)}}$ (44)
$=\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{1}\leqslant\tau_{2},t_{1}\leqslant\tau_{3},t_{1}\leqslant\tau_{4}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{1}}\mathbf{z}_{t_{1}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{1}}\right]$
@@
$\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{3}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{1}\leqslant\tau_{2},t_{3}\leqslant\tau_{3},t_{3}\leqslant\tau_{4}}\bm{1}_{t_{1}\neq t_{3}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{1}}\mathbf{z}_{t_{3}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{3}}\right]$
@@
$\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{1}\leqslant\tau_{3},t_{2}\leqslant\tau_{4}}\bm{1}_{t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{1}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{2}}\right]$
@@
${\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{2}\leqslant\tau_{3},t_{1}\leqslant\tau_{4}}\bm{1}_{t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}_{t_{1}}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}_{t_{2}}\mathbf{z}_{t_{2}}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}_{t_{1}}\right]}$ (45)
$=\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{1}\leqslant\tau_{2},t_{1}\leqslant\tau_{3},t_{1}\leqslant\tau_{4}}\operatorname*{\mathbb{E}}\left[\mathbf{z}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}\mathbf{z}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}\right]$
@@
$\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{3}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{1}\leqslant\tau_{2},t_{3}\leqslant\tau_{3},t_{3}\leqslant\tau_{4}}\bm{1}_{t_{1}\neq t_{3}}\operatorname*{\mathbb{E}}\left[\mathbf{z}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}\right]\operatorname*{\mathbb{E}}\left[{\mathbf{z}^{\prime}}^{\top}A(\tau_{3},\tau_{4}){\mathbf{z}^{\prime}}\right]$
@@
$\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{1}\leqslant\tau_{3},t_{2}\leqslant\tau_{4}}\bm{1}_{t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}^{\prime}\mathbf{z}^{\top}A(\tau_{3},\tau_{4})\mathbf{z}^{\prime}\right]$
@@
${\phantom{=}+\sum_{\tau_{1},\tau_{2},\tau_{3},\tau_{4}=t_{0}}^{t}\frac{1}{\tau_{1}\tau_{2}\tau_{3}\tau_{4}}\sum_{t_{1},t_{2}}\bm{1}_{t_{1}\leqslant\tau_{1},t_{2}\leqslant\tau_{2},t_{2}\leqslant\tau_{3},t_{1}\leqslant\tau_{4}}\bm{1}_{t_{1}\neq t_{2}}\operatorname*{\mathbb{E}}\left[\mathbf{z}^{\top}A(\tau_{1},\tau_{2})\mathbf{z}^{\prime}\mathbf{z}^{\prime\top}A(\tau_{3},\tau_{4})\mathbf{z}\right]}$ (46)
@@
As $W$ is symmetric and stochastic, then the module of its largest eigenvalue is equal to $1$. The vector $\mathbf{1}_{n}$ is both a left and right eigenvector of $W$ relative to the simple eigenvalue $1$. Then, $P$ is the projector onto the null space of $W-I$ along the range of $W-I$ (Meyer 2023)[p.518]. The spectral theorem leads us to conclude that the eigenvalues of $W-P$ (counted with their multiplicity) are then all eigenvalues of $W$ except $1$ and with the addition of 00. We can then conclude that $\lVert W-P\rVert=\rho(W-P)=\lambda_{2}(W)$. $W$ is irreducible with non-negative elements on the diagonal, then it is primitive (Meyer 2023)[Example 8.3.3], i.e., $1$ is the only eigenvalue on the unit circle.
@@
where (57) follows from Lemma A.1 in (Shalev-Shwartz and Ben-David 2014). For $a<1$, it is easy to check that for $x\geqslant 1$, $x\geqslant\frac{a}{4}\ln^{2}x$ holds unconditionally. ∎
F.3 Bounding the $4$-th Moment of the Estimation Error.
@@
F.4 $(\varepsilon,\delta)$-Bounds: Proof of Theorem 3